A 40-kilogram mass is moving across a horizontal surface at 5.0 m/s. What is the magnitude of the net force required to bring the mass to a stop in 8.0 seconds? 


1
Expert's answer
2022-01-11T04:13:10-0500

Given:

"m=\rm 40\: kg"

"v_i=\rm 5.0\: m/s"

"v_f=\rm 0\: m/s"

"\Delta t=\rm 8\: s"


The Newton's second law says

"F=ma=m\frac{\Delta v}{\Delta t}"

"F=\rm 40*\frac{5.0-0}{8}=25\: N"